# Structure-Informed Estimation for Pilot-Limited MIMO Channels via Tensor Decomposition

### v1

@@ 5.1 Setup
For synthetic experiments, the wideband channel is represented as a third-order tensor $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}$ with $N_{r}=N_{t}=32$ antennas and $N_{f}=128$ subcarriers, generated according to the multipath model in (7) with $L$ dominant paths. Angles satisfy $\theta_{\ell},\phi_{\ell}\sim\mathcal{U}[-\pi/2,\pi/2]$ (angular spread $180^{\circ}$), complex gains $\alpha_{\ell}$ are drawn i.i.d. from $\mathcal{CN}(0,1)$ and normalized to unit total power, and normalized delays satisfy $\tau_{\ell}\sim\mathcal{U}[0,1)$. Uniform linear arrays (ULA) with half-wavelength spacing are used at both ends [2, 7]. Pilot acquisition is modeled as entry-wise random sampling over an index set $\Omega$ with pilot ratio $|\Omega|/(N_{r}N_{t}N_{f})$ [33, 26]. Additive white Gaussian noise (AWGN) is applied only to observed entries, and SNR is defined with respect to the average power of observed coefficients.
The following estimators are compared:
  • LS (zero-fill): baseline estimator setting unobserved entries to zero;
  • OMP: orthogonal matching pursuit with Kronecker ULA steering dictionary (Rx/Tx), applied independently per subcarrier [7];
@@
  • CP: rank-$R$ CP decomposition using Tucker completion as low-rank proxy, with data-consistency projection [23];
  • Tensor-NN: proposed hybrid architecture combining Tucker completion with lightweight 3D U-Net for residual learning.
Performance is measured by normalized mean-squared error (NMSE),
@@
averaged over Monte Carlo (MC) realizations. Table 2 summarizes experiment settings.
Table 2: Simulation parameters.
Parameter Exp. 1 Exp. 2 Exp. 3
Sweep variable Pilot ratio SNR Pilot ratio
Pilot ratio 2–20% 8% 2–20%
SNR (dB) 10 5-5 to 20 10
$L$ paths 5 5 5
$N_{r}\times N_{t}\times N_{f}$ $\times 32\times 128 32×32$$\times 32\times 128$ $16\times 32\times 64$
Bandwidth 50 MHz
Tucker ranks $(4,4,6)$ $(5,5,6)$ $(4,4,8)$
@@
Dataset Synthetic Synthetic DeepMIMO
MC runs 500 500 500
5.1.1 DeepMIMO Dataset
To assess the proposed framework under realistic propagation conditions, ray-tracing-based channels are taken from the DeepMIMO dataset [34]. Experiments use the ASU_Campus_3p5 scenario (DeepMIMO v4 pre-release), generated with the Remcom Wireless InSite ray tracer. This scenario models an outdoor campus environment at 3.5 GHz with dense building geometry and heterogeneous scattering objects, providing a substantially more challenging testbed than idealized sparse multipath models. Detailed ray-tracing parameters (path depth, reflections, diffractions) are defined by the DeepMIMO dataset configuration and documented in [34].
Point-to-point MIMO-OFDM channels are constructed using ULAs with $N_{r}=16$ receive antennas, $N_{t}=32$ transmit antennas, and $N_{f}=64$ subcarriers over 50 MHz bandwidth. Channels are taken directly from DeepMIMO without path truncation and normalized to unit average power; low-rank structure is therefore empirical rather than enforced by construction.
Training and test sets are defined by random subsampling of user positions with fixed seeds for reproducibility. This partitioning does not enforce explicit geographic separation between training and test locations; evaluating spatial generalization under strict geographic partitioning remains a direction for future work.
@@ 5.2 Experiment 1: Tensor Completion under Pilot Scarcity
Fig. 2 evaluates NMSE as a function of pilot ratio at SNR =10=10 dB with $L=5$ paths. This experiment highlights the central contribution: tensor completion yields meaningful reconstruction gains under severe pilot undersampling.
LS yields NMSE near unity across all pilot ratios, confirming that naive zero-filling cannot recover channel structure under sparse sampling. OMP achieves moderate gains at higher pilot ratios but exhibits a performance floor due to dictionary mismatch and per-subcarrier processing that does not exploit frequency-domain correlations.
In contrast, Tucker completion exploits low-rank structure across all three tensor modes and achieves substantial gains even at very low pilot densities (4–6%), with NMSE improving monotonically as more pilots become available. At 10% pilots, Tucker attains approximately one order of magnitude improvement over LS. These results demonstrate that algebraic tensor priors effectively regularize the ill-posed inverse problem arising from pilot-limited observation.
Figure 2: NMSE versus pilot ratio for synthetic channels (SNR =10=10 dB, $L=5$ paths, Tucker ranks $(4,4,6)$, 500 MC runs). Tucker completion improves over LS by $\sim$2–12 dB across the sweep, exceeding 10 dB for pilot ratios $\geq 0.10$, and it outperforms OMP at all pilot ratios.
@@ 5.3 Experiment 2: CP versus Tucker Decomposition
Fig. 3 compares Tucker and CP as a function of SNR at 8% pilot ratio. This experiment addresses the practical question of decomposition selection based on channel characteristics.
Both tensor methods substantially outperform LS and OMP across all SNR values. At low SNR (5-5 and 0 dB), the two methods are comparable, with CP slightly worse (sub‑1 dB difference). As SNR increases, CP progressively outperforms Tucker; at SNR =20=20 dB, CP achieves about a 3.6 dB NMSE advantage.
This behavior is consistent with the synthetic channel model in (7), which is a sum of $L$ rank‑one components. CP provides a tighter parameterization when the channel matches this specular multipath structure, whereas Tucker introduces additional degrees of freedom through its core tensor. Under model mismatch (e.g., richer scattering or non‑ideal effects, as in DeepMIMO), Tucker’s flexibility can become advantageous.
Figure 3: Tucker versus CP as a function of SNR ($\rho=8\%$, $L=5$ paths, Tucker ranks $(5,5,6)$, CP rank $R=5$, 500 MC runs). CP performs better than Tucker at moderate-to-high SNR for specular channels consistent with the multipath model.
@@ 5.4 Experiment 3: Hybrid Tensor–Neural Estimation on DeepMIMO
Fig. 4 evaluates the complete framework on ray-tracing-based DeepMIMO channels at SNR =10=10 dB. These channels exhibit higher effective rank due to diffuse scattering and environment-specific propagation characteristics not captured by the idealized model in (7).
Pure tensor methods (Tucker and CP) continue to outperform LS on DeepMIMO channels, but the performance gap relative to synthetic channels narrows due to model mismatch. A purely data-driven ResNet baseline (convolutional neural network (CNN)) is also included [35, 36, 37]. The ResNet is a 3D architecture operating on the real/imaginary channels of the observation and the pilot mask (3-channel input), with a $3\!\times\!3\!\times\!3$ Conv3D stem, followed by 4 residual blocks (each with two $3\!\times\!3\!\times\!3$ Conv3D layers and ReLU), and a final $3\!\times\!3\!\times\!3$ Conv3D to predict the 2-channel complex output (base channels =32=32).
Tensor-NN consistently outperforms pure tensor methods across all pilot ratios. At 8% pilots, Tensor-NN achieves NMSE $\approx 0.072$ compared to $\approx 0.128$ (Tucker) and $\approx 0.131$ (CP), representing approximately 44% improvement; the ResNet baseline yields NMSE $\approx 0.088$ at the same pilot density. Tensor-NN overtakes ResNet from $\rho\geq 0.04$ and maintains a clear gap thereafter. At 20% pilots, Tensor-NN achieves NMSE $\approx 0.028$ versus $\approx 0.037$ (Tucker), a 24% gain, while ResNet attains $\approx 0.054$.
ResNet training methodology
The ResNet baseline is trained offline on paired $(\mathcal{H},\mathcal{Y}_{\Omega})$ samples. Inputs are formed by stacking the real and imaginary parts of the noisy pilot observations together with the binary pilot mask (3-channel input), and the network outputs the real/imaginary channels of $\mathcal{H}$. Training minimizes the mean‑squared error between the predicted and true channel tensors, using Adam optimization and early stopping. To promote robustness, pilot ratios and SNR are sampled from ranges during training (multi‑regime training), and random pilot masks are generated per sample. In the pure ResNet setting, no explicit pilot‑consistency term is used (i.e., $\lambda_{\mathrm{pilot}}=0$). For the DeepMIMO experiment, training uses 1500 samples for training and 300 for validation, 25 epochs, base channels =32=32, and 4 residual blocks.
Figure 4: Performance on DeepMIMO ray-tracing channels (ASU_Campus_3p5 scenario, SNR =10=10 dB) versus pilot ratio. A purely data-driven ResNet baseline is included for reference. Tensor-NN outperforms Tucker and CP and surpasses ResNet for pilot ratios $\rho\geq 0.04$.
@@ 5.5 Experiment 4: Sample Complexity Analysis
This experiment empirically characterizes the sample complexity of tensor completion by determining the recovery threshold of Tucker reconstruction as a function of the number of observed entries $|\Omega|$. Synthetic channels are generated according to the model in Section 5.1, with $N_{r}=N_{t}=32$ antennas, $N_{f}=128$ subcarriers, and $L\in\{2,3,5,8,10,15\}$ propagation paths. Pilot ratios are swept from 1% to 25% across 14 sample points, with finer spacing at low pilot ratios (and consequently finer resolution near the transition region). Results are averaged over 100 Monte Carlo realizations, and successful recovery is defined by $\mathrm{NMSE}\leq 10^{-2}$, which determines the empirical threshold $|\Omega|_{\min}$. Following the implementation, Tucker ranks are adapted to $L$ via $(R_{r},R_{t},R_{f})=(L,L,L+1)$ (capped by $N_{f}$).
@@
Fig. 5 plots NMSE versus pilot ratio $\rho$ for each value of $L$ at $\mathrm{SNR}\in\{10,20,30\}$ dB. Two consistent trends emerge. First, the recovery transition shifts toward higher pilot densities as $L$ increases, reflecting the growing intrinsic dimensionality of the channel model. Second, higher SNR sharpens the transition and lowers the NMSE once the recovery region is reached. At $\mathrm{SNR}=10$ dB, no configuration achieves $\mathrm{NMSE}\leq 10^{-2}$ within the tested 25% pilot range, indicating a noise-limited regime where increasing pilot density alone cannot compensate for measurement noise. The curves remain monotonic but do not exhibit a clear threshold within the tested range.
@@
Table 3 reports the minimum pilot ratio $\rho_{\min}$ and corresponding observation count $|\Omega|_{\min}$ required to achieve $\mathrm{NMSE}\leq 10^{-2}$. At $\mathrm{SNR}=10$ dB, no configuration reaches this threshold within the tested range. For $\mathrm{SNR}\in\{20,30\}$ dB, thresholds lie in the 12–25% pilot range and increase monotonically with $L$, confirming that sample complexity is dominated by intrinsic channel structure rather than noise level.
@@
20 2 0.120 15728.6 40.96
20 3 0.150 19660.8 34.13
20 5 0.150 19660.8 20.48
20 8 0.200 26214.4 17.07
20 10 0.200 26214.4 13.65
20 15 0.250 32768.0 11.38
30 2 0.120 15728.6 40.96
30 3 0.120 15728.6 27.31
30 5 0.150 19660.8 20.48
30 8 0.150 19660.8 12.80
30 10 0.200 26214.4 13.65
30 15 0.250 32768.0 11.38
To interpret these results, we compare against the heuristic degrees-of-freedom scaling for low-rank tensor recovery:
@@
where $C$ denotes the oversampling factor. The rightmost column of Table 3 lists the empirically observed $C$ for each configuration. The measured values ($C\approx 11$4141) exceed theoretical minima established for incoherent low-rank tensors under idealized conditions [38]. This gap is attributable to three factors: (i) Tucker decomposition introduces additional core tensor parameters absent in CP; (ii) finite-sample effects and interpolation in threshold estimation; and (iii) suboptimality of the alternating projection algorithm under noisy, sparse observations.
We emphasize that the scaling in (18) is CP-inspired and serves here as a qualitative proxy for interpreting recovery thresholds. For Tucker models, the exact degrees of freedom additionally depend on the multilinear ranks $(R_{r},R_{t},R_{f})$ and core tensor size. Nonetheless, the empirical results demonstrate that the minimum observation count scales predominantly with intrinsic channel complexity—captured by the number of dominant propagation paths and effective subspace dimensions—rather than with the ambient tensor size $N_{r}N_{t}N_{f}$.
Fig. 7 compares empirical thresholds against the heuristic bound (18) with $C=5$. At $\mathrm{SNR}=10$ dB, no empirical thresholds are available within the tested range, so only the bound is shown. Despite the constant-factor gap, empirical $|\Omega|_{\min}$ scales approximately linearly with both $L$ and $(N_{r}+N_{t}+N_{f})$, confirming that sample complexity is governed by intrinsic model dimensionality rather than ambient tensor size.
(a) SNR = 10 dB
(b) SNR = 20 dB
(c) SNR = 30 dB
Figure 7: Empirical $|\Omega|_{\min}$ versus the heuristic bound in (18) with $C=5$ for $\mathrm{SNR}\in\{10,20,30\}$ dB.
@@
5.6 Computational Complexity
Table 4 summarizes the computational complexity and measured execution times for each method. The dominant cost in tensor-based approaches is the low-rank projection step. For Tucker completion via truncated HOSVD, each iteration requires $\mathcal{O}(N_{r}N_{t}N_{f}\cdot R_{\max})$ operations, where $R_{\max}=\max(R_{r},R_{t},R_{f})$ denotes the largest multilinear rank. The algorithm converges in 20 iterations for synthetic channels and 30 iterations for DeepMIMO experiments, reflecting the higher effective rank of ray-tracing-based channels.
CP employs Tucker completion as a low-rank proxy followed by data-consistency enforcement, yielding comparable per-iteration cost. The hybrid Tensor-NN method adds a single forward pass through the 3D U-Net (${\sim}50$k parameters), contributing marginal overhead relative to the tensor completion stage.
All methods achieve execution times below 40 ms per channel realization on an Apple M3 Pro processor, well within the coherence time of typical wireless channels (1–10 ms at pedestrian speeds, 0.1–1 ms at vehicular speeds). The modest overhead of Tensor-NN (${\sim}20$ ms beyond Tucker) is justified by the 24–44% NMSE improvement observed on realistic channels. For deployment on resource-constrained hardware, further acceleration is possible through GPU batching, quantization, or algorithmic simplifications such as reduced iteration count with warm-start initialization.
Table 4: Computational complexity and execution time per channel realization ($N_{r}=N_{t}=32$, $N_{f}=128$, pilot ratio $8\%$, Apple M3 Pro).
Method Complexity Iter. Time
LS (zero-fill) $\mathcal{O}(|\Omega|)$ <1<1 ms
Tucker $\mathcal{O}(N_{r}N_{t}N_{f}\cdot R_{\max})$ 20 $\approx 15$ ms
CP $\approx$ Tucker 30 $\approx 25$ ms
Tensor-NN Tucker + $\mathcal{O}(N_{r}N_{t}N_{f}\cdot C^{2})$ 20+1 $\approx 35$ ms

### v2

@@ 5.1 Setup
For synthetic experiments, the setup represents the wideband channel as a third-order tensor $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}$ with $N_{r}=N_{t}=32$ antennas and $N_{f}=128$ subcarriers, generating it according to the multipath model in (7) with $L$ dominant paths. Angles satisfy $\theta_{\ell},\phi_{\ell}\sim\mathcal{U}[-\pi/2,\pi/2]$ (angular spread $180^{\circ}$), complex gains $\alpha_{\ell}$ are drawn i.i.d. from $\mathcal{CN}(0,1)$ and normalized to unit total power, and normalized delays satisfy $\tau_{\ell}\sim\mathcal{U}[0,1)$. ULA with half-wavelength spacing serve at both ends [2, 7]. The setup models pilot acquisition as entry-wise random sampling over an index set $\Omega$ with pilot ratio $\rho=|\Omega|/(N_{r}N_{t}N_{f})$ [33, 26]. The model applies AWGN only to observed entries. The noise variance is set according to a global SNR reference,
$\sigma_{n}^{2}=\frac{\|\mathcal{H}\|_{F}^{2}/(N_{r}N_{t}N_{f})}{10^{\mathrm{SNR}/10}},$ (16)
where $\|\mathcal{H}\|_{F}^{2}/(N_{r}N_{t}N_{f})$ denotes the average power per channel coefficient over the full tensor, independent of the pilot set $\Omega$ and pilot ratio $\rho$. This definition is applied identically across all experiments, including the DeepMIMO scenario (Experiment 3), ensuring strict SNR comparability across pilot densities.
For synthetic experiments, the channel model operates in a normalized discrete-frequency domain, where delays are specified in sample units, leaving bandwidth undefined. Since the evaluation targets reconstruction accuracy rather than physical delay interpretation, results remain invariant to bandwidth scaling.
The following estimators enter the comparison:
  • LS: baseline estimator setting unobserved entries to zero;
  • LMMSE: linear MMSE estimator with covariance estimated from channel samples, used as a practical statistical reference baseline;
  • orthogonal matching pursuit (OMP): joint sparse recovery using simultaneous OMP (SOMP) with shared support across subcarriers and a Kronecker ULA steering dictionary (Rx/Tx). For notation consistency, this joint SOMP curve carries the label “OMP” in the plots [7];
@@
  • CP: rank-$R$ CP completion with missing-data-aware WALS, selecting the best restart by observed-data fit [23];
  • Tensor-NN: proposed hybrid architecture combining low-rank tensor completion and a lightweight 3D U-Net for residual learning.
The experiments measure performance by normalized mean-squared error (NMSE),
@@
averaging each result over Monte Carlo (MC) realizations. Table 2 summarizes experiment settings.
Table 2: Simulation parameters. Exp. 4 employs 100 Monte Carlo realizations, as its objective is recovery-threshold (sample-complexity) characterization rather than statistical performance averaging.
Parameter Exp. 1 Exp. 2 Exp. 3 Exp. 4
Sweep variable Pilot ratio SNR Pilot ratio Pilot ratio / $L$ paths
Pilot ratio 2–20% 8% 2–20% 1–25%
SNR (dB) 10 5-5 to 20 10 10
$L$ paths 5 5 5 $\{2,3,5,8,10,15\}$
$N_{r}\times N_{t}\times N_{f}$ $\times 32\times 128 32×32$$\times 32\times 128$ $16\times 32\times 64$ $32\times 32\times 128$
Tucker ranks $(4,4,6)$ $(5,5,6)$ $(4,4,8)$ $(L,L,L+1)$
@@
Dataset Synthetic Synthetic DeepMIMO Synthetic
MC runs 500 500 500 100
Experiment 1 uses ranks $(4,4,6)$ to assess robustness under mild rank underestimation; Experiment 2 uses $(5,5,6)$ to match the channel rank more closely and isolate the CP–Tucker comparison at higher SNR.
This work also evaluates realistic propagation conditions using ray-tracing-based channels from the DeepMIMO dataset [34]. The experiments adopt the ASU_Campus_3p5 scenario (DeepMIMO v4 pre-release), generated with Remcom Wireless InSite at 3.5 GHz.
Experiment 3 employs ULA with $N_{r}=16$ receive antennas and $N_{t}=32$ transmit antennas, and OFDM with $N_{f}=64$ subcarriers over 50 MHz bandwidth, yielding $\mathcal{H}\in\mathbb{C}^{16\times 32\times 64}$ (Table 2). The study uses channels without path truncation and normalizes them to unit average power; low-rank structure therefore emerges empirically. The protocol forms training and test sets via random subsampling of user locations with fixed seeds for reproducibility. The partitioning enforces no explicit geographic separation; strict spatial generalization remains future work. Channels are normalized to unit average power prior to noise addition, so that $\|\mathcal{H}\|_{F}^{2}/(N_{r}N_{t}N_{f})=1$ and the noise variance reduces to $\sigma_{n}^{2}=10^{-\mathrm{SNR}/10}$, consistently with the global SNR definition in (16).
Experiment 4 adopts the same synthetic channel model with $N_{r}=N_{t}=32$ and $N_{f}=128$, varying the number of paths $L\in\{2,3,5,8,10,15\}$ and sweeping pilot ratios from $1\%$ to $25\%$ across 14 sample points, with Tucker ranks set to $(L,L,L+1)$. Each result averages over 100 Monte Carlo realizations, and the analysis defines successful recovery as $\mathrm{NMSE}\leq 10^{-2}$.
@@ 5.2 Experiment 1: Tensor Completion under Pilot Scarcity
Fig. 4 shows NMSE as a function of pilot ratio at SNR =10=10 dB with $L=5$ paths (500 MC runs, global-SNR setting). This experiment highlights the benefit of Tucker tensor completion under pilot undersampling. CP is excluded from this comparison as its performance relative to Tucker under the synthetic specular channel model is isolated in Experiment 2 (Section 5.3), which sweeps SNR at fixed pilot ratio and provides a cleaner characterization of the CP–Tucker trade-off; furthermore, CP exhibits heavy-tail divergence at low pilot densities (as confirmed in Experiment 3), which would compress the vertical scale and obscure the Tucker–baseline comparison that is the focus of this experiment.
LS remains weak across the sweep, ranging from approximately 0.08-0.08 dB at $\rho=2\%$ to 0.86-0.86 dB at $\rho=20\%$, confirming that naive interpolation cannot recover channel structure from sparse pilots. LMMSE provides only modest improvement, decreasing from about 0.16-0.16 dB to 1.28-1.28 dB over the same range. OMP improves over LS/LMMSE and reaches around 3.5-3.5 dB at higher pilot ratios, but still exhibits a clear performance floor.
Tucker completion improves monotonically with pilot density, from about 1.80-1.80 dB at $\rho=2\%$ to 12.79-12.79 dB at $\rho=20\%$. At $\rho=10\%$, Tucker achieves $\mathrm{NMSE}\approx-11.25$ dB, corresponding to gains of $10.88$ dB over LS ($\approx-0.37$ dB), $10.52$ dB over LMMSE, and $7.83$ dB over OMP ($\approx-3.42$ dB). Notably, at the most extreme sparsity point ($\rho=2\%$), OMP is slightly better than Tucker (about $0.67$ dB), but Tucker becomes clearly dominant from $\rho\geq 4\%$ onward.
Figure 4: NMSE versus pilot ratio for synthetic specular channels (SNR =10=10 dB, $L=5$, Tucker ranks $(4,4,6)$, 500 MC runs).
@@ 5.3 Experiment 2: CP versus Tucker Decomposition
Fig. 5 compares Tucker and CP as a function of SNR at $\rho=8\%$. This experiment addresses decomposition choice under the synthetic specular channel model.
Both tensor methods outperform LS and LMMSE across the full SNR range, and CP consistently outperforms Tucker. At low SNR (5-5 dB), CP already provides a slight gain over Tucker (CP $\approx-1.82$ dB versus Tucker $\approx-1.37$ dB, about $0.45$ dB). The gap widens markedly with SNR: at $20$ dB, CP reaches $\mathrm{NMSE}\approx-25.00$ dB versus Tucker $\approx-11.89$ dB, i.e., about $13.11$ dB advantage. Compared to OMP, Tucker is slightly worse at 5-5 dB (OMP $\approx-2.09$ dB) but becomes clearly better from 0 dB onward.
This behavior is consistent with the synthetic model in (7), where the channel consists of a small number of rank-one components. In this regime, CP provides a tighter representation than Tucker’s full-core parameterization.
Figure 5: NMSE versus SNR for synthetic specular channels ($\rho=8\%$, $L=5$, Tucker ranks $(5,5,6)$, CP rank $R=5$, 500 MC runs).
@@ 5.4 Experiment 3: Hybrid Tensor–Neural Estimation on DeepMIMO
Fig. 6 evaluates the complete framework on ray-tracing-based DeepMIMO channels at SNR =10=10 dB. These channels exhibit higher effective rank and stronger model mismatch relative to the idealized synthetic setting.
The comparison includes LS, LMMSE, OMP, Tucker, CP, Tensor-NN, and a purely data-driven ResNet baseline. Tensor-NN achieves the best overall performance across all pilot ratios, with NMSE decreasing from approximately 4.56-4.56 dB at $\rho=2\%$ to approximately 18.43-18.43 dB at $\rho=20\%$.
At $\rho=8\%$, Tensor-NN reaches approximately 13.72-13.72 dB, improving by about $2.26$ dB over CP ($\approx-11.46$ dB), $4.80$ dB over Tucker ($\approx-8.93$ dB), and $5.85$ dB over LMMSE ($\approx-7.87$ dB). At $\rho=20\%$, Tensor-NN remains best ($\approx-18.43$ dB), with a positive margin over CP ($\approx-17.43$ dB, about $1.00$ dB advantage).
CP shows severe instability at extreme pilot scarcity ($\rho=2\%$), with heavy-tailed outliers in Monte Carlo runs: its mean NMSE becomes strongly degraded (about +13.96+13.96 dB), consistent with rare divergent cases dominating the average.
Comparison with the ResNet baseline
A purely data-driven ResNet baseline [35, 36, 37] is included to quantify the contribution of tensor structural priors over unconstrained deep learning. The ResNet is a 3D architecture operating on the real and imaginary parts of the observation and the pilot mask (3-channel input), with a $3\!\times\!3\!\times\!3$ Conv3D stem followed by 4 residual blocks (each comprising two $3\!\times\!3\!\times\!3$ Conv3D layers with ReLU activations) and a final $3\!\times\!3\!\times\!3$ Conv3D predicting the 2-channel complex output (base channels =32=32). Training uses paired $(\mathcal{H},\mathcal{Y}_{\Omega})$ samples optimized with Adam under multi-SNR and multi-pilot regimes, with pilot-consistency weight $\lambda_{\mathrm{pilot}}=0.1$.
Tensor-NN outperforms ResNet at all pilot ratios: by $1.53$ dB at $\rho=2\%$, $3.64$ dB at $\rho=4\%$, $3.58$ dB at $\rho=6\%$, $3.67$ dB at $\rho=8\%$, $3.91$ dB at $\rho=10\%$, $4.91$ dB at $\rho=15\%$, and $5.67$ dB at $\rho=20\%$.
Overall, the hybrid Tensor-NN framework provides the most robust reconstruction across the full pilot-density range, combining the structural efficiency of tensor decomposition with the flexibility of neural residual correction under realistic DeepMIMO propagation conditions.
Figure 6: NMSE versus pilot ratio (DeepMIMO ASU_Campus_3p5, SNR =10=10 dB, 500 MC runs).
@@ 5.5 Experiment 4: Sample Complexity Analysis
This experiment empirically characterizes the sample complexity of tensor completion by determining the recovery threshold of Tucker reconstruction as a function of the number of observed entries $|\Omega|$. The setup generates synthetic channels according to the model in Section 5.1, with $N_{r}=N_{t}=32$ antennas, $N_{f}=128$ subcarriers, and $L\in\{2,3,5,8,10,15\}$ propagation paths. The sweep covers pilot ratios from 1% to 25% across 14 sample points, with finer spacing at low pilot ratios (and consequently finer resolution near the transition region). Each result averages over 100 Monte Carlo realizations, and the analysis defines successful recovery as $\mathrm{NMSE}\leq 10^{-2}$, which determines the empirical threshold $|\Omega|_{\min}$. Following the implementation, the experiments adapt Tucker ranks to $L$ via $(R_{r},R_{t},R_{f})=(L,L,L+1)$ (capped by $N_{f}$).
@@
Fig. 7 plots NMSE versus pilot ratio $\rho$ for each value of $L$ at $\mathrm{SNR}\in\{10,20,30\}$ dB. Two consistent trends emerge. First, the recovery transition shifts toward higher pilot densities as $L$ increases, reflecting the growing intrinsic dimensionality of the channel model. Second, higher SNR sharpens the transition and lowers NMSE once the recovery region is reached. At $\mathrm{SNR}=10$ dB, no configuration achieves $\mathrm{NMSE}\leq 10^{-2}$ within the tested 25% pilot range, indicating a noise-limited regime where increasing pilot density alone cannot compensate for measurement noise. The curves remain monotonic but exhibit no clear threshold within the tested range.
@@
Table 3 reports the minimum pilot ratio $\rho_{\min}$ and corresponding observation count $|\Omega|_{\min}$ required to achieve $\mathrm{NMSE}\leq 10^{-2}$. At $\mathrm{SNR}=10$ dB, no configuration reaches this threshold within the tested pilot range. For $\mathrm{SNR}\in\{20,30\}$ dB, the thresholds fall within the 12–25% pilot interval and exhibit an overall increasing trend with $L$, indicating that sample complexity is primarily governed by intrinsic channel structure rather than noise level.
@@
20 2 0.120 15,729 40.96
20 3 0.150 19,661 34.13
20 5 0.150 19,661 20.48
20 8 0.200 26,214 17.07
20 10 0.200 26,214 13.65
20 15 0.250 32,768 11.38
30 2 0.120 15,729 40.96
30 3 0.120 15,729 27.31
30 5 0.150 19,661 20.48
30 8 0.150 19,661 12.80
30 10 0.200 26,214 13.65
30 15 0.250 32,768 11.38
To interpret these results, this article compares them against the heuristic degrees-of-freedom scaling commonly used in low-rank tensor recovery:
@@
where $C$ denotes an oversampling factor capturing deviations from the idealized identifiability limit. This expression is CP-inspired: it counts $L$ factor vectors of lengths $N_{r}$, $N_{t}$, and $N_{f}$, totalling $L(N_{r}+N_{t}+N_{f})$ free parameters.
For Tucker decompositions with multilinear ranks $(R_{r},R_{t},R_{f})$, the effective degrees of freedom are
$d_{\text{Tucker}}=R_{r}R_{t}R_{f}+R_{r}N_{r}+R_{t}N_{t}+R_{f}N_{f},$ (19)
comprising the core tensor ($R_{r}R_{t}R_{f}$ entries) and the three factor matrices ($R_{n}N_{n}$ entries each). For the ranks $(R_{r},R_{t},R_{f})=(L,L,L+1)$ used in Experiment 4,  (19) yields values ranging from $d_{\text{Tucker}}=524$ at $L=2$ to $d_{\text{Tucker}}=6608$ at $L=15$, compared with $L(N_{r}+N_{t}+N_{f})\in\{384,\,2880\}$ from (18). The ratio $d_{\text{Tucker}}/[L(N_{r}+N_{t}+N_{f})]$ grows from $\approx\!1.36$ at $L=2$ to $\approx\!2.29$ at $L=15$, reflecting the cubic growth of the core-tensor term $R_{r}R_{t}R_{f}\sim L^{3}$ relative to the linear scaling of the CP heuristic.
The rightmost column of Table 3 lists the empirically observed values of $C$. The measured range ($C\approx 11$$41$) exceeds theoretical minima established for incoherent low-rank tensors under noiseless and asymptotic conditions [38]. Several practical factors contribute to this gap. First, the CP-inspired reference (18) increasingly underestimates true Tucker complexity as $L$ grows, since it does not account for the core-tensor degrees of freedom in  (19); the empirical $C$ must therefore absorb this structural gap in addition to other sources of inefficiency. Second, finite-sample effects and interpolation in threshold estimation introduce bias. Third, practical convergence limitations of the alternating projection algorithm under noisy and sparse observations may prevent attainment of the optimal recovery boundary.
Nonetheless, the empirical evidence indicates that the minimum observation count scales predominantly with intrinsic channel complexity—characterized by the number of dominant propagation paths and effective subspace dimensions—rather than with the ambient tensor size $N_{r}N_{t}N_{f}$, consistent with Table 3.